# Child Speech Recognition in Human-Robot Interaction: Problem Solved?

### v1

@@ 3.4. Energy consumption
As concerns have been raised over the energy consumption and consequently carbon emissions of state-of-the-art machine learning (Lacoste et al., 2019), we think it is valuable to consider these for ASR systems. We only have been able to measure Whisper’s consumption. Per hour of transcribed data, the largest and best-performing model (large-v3) consumes $32.3Wh$ and produces $7.7gCO_{2}eq$.
4. Conclusion
@@
Based on our evaluation, we can make the following recommendations, updating or overriding those made in (Kennedy et al., 2017):
Recognition performance.:
@@
The recognition performance has improved dramatically for state-of-the-art ASR, with the best models of 2024 showing over 60% fewer transcription errors than in 2017. Still, adult-like recognition is not available yet, but the semantic content of children’s speech is now sufficiently transcribed to offer potential for robust spoken interaction, especially if other components within the dialogue management –such as large language models– can cover for suboptimal or even failing ASR.
Responsiveness.:
@@
Acknowledgements.
This research received funding from imec (Smart Education), the Flemish Government (AI Research Program) and the Horizon Europe VALAWAI project (grant agreement number 101070930). We are indebted to the authors of (Kennedy et al., 2016) and (Kennedy et al., 2017) for making the recordings and transcriptions available.

### v2

@@ 3.6 Energy consumption
As concerns have been raised over the energy consumption and consequently carbon emissions of state-of-the-art machine learning [6], we think it is valuable to consider these for ASR systems. We only have been able to measure Whisper’s consumption. Per hour of transcribed data, the largest and best-performing model (large-v3) consumes $32.3Wh$ and produces $7.7gCO_{2}eq$. Note that this energy consumption refers to the transcription of many samples back-to-back, while real-time interactive systems would not be able to work as efficiently.
4 Conclusion
@@
Based on our evaluation, we can make the following recommendations, updating or overriding those made in [3]:
Recognition performance.
@@
The recognition performance has improved dramatically for state-of-the-art ASR, with the best models of 2024 showing over 60% fewer transcription errors than in 2017. However, performance falters when input samples become much shorter or longer than a single sentence. Model priming is able to improve performance, especially in very structured settings, e.g. when a number of fixed options are expected. Adult-like recognition is not available yet, but the semantic content of children’s speech is now sufficiently transcribed to offer potential for robust spoken interaction. Small errors can often be caught by other components of the dialogue management system —such as large language models— resulting in a higher task accuracy than the reported accuracy of ASR systems.
Responsiveness.
@@
{credits}
4.0.1 Acknowledgements
This research received funding from imec (Smart Education), the Flemish Government (AI Research Program) and the Horizon Europe VALAWAI project (grant agreement number 101070930). We are indebted to the authors of [4] and [3] for making the recordings and transcriptions available.